Your Parents are planning to help you while you are in college. He has vowed that he will give you $170 each month for all 4 years while you are in college. Given an interest rate of .22 percent per month, what is the present value of these cash flows worth to you when you first start college? Multiple Choice $7,735.86 $7,911.24 $7,486.43 $7,553.03 $7,216.86
Answer #1

Given that,

Monthly deposit, PMT = $170

years of deposit, t = 4

Total monthly payments made N = 4*12 = 48

Monthly rate r = 0.22%

So, present value of these cash flow can be calculated using PV formula of annuity,

PV = PMT*(1 - (1+r)^(-N))/r = 170*(1 - 1.0022^(-48))/0.0022 = $7735.86

So, Present value of these cash flows = $7735.86

Option A is correct.